A sample of 0.892 g of potassium chloride (KCl) is dissolved in 54.6 g of water. What is the percent by mass of KCl in this solution?
Calculate the percent by volume of 20 mL of alcohol diluted in water to produce 250mL solution.
1
Expert's answer
2020-10-23T07:17:53-0400


The percent by mass of KCl in this solution is


"\omega=\frac{m_\text{KCl}}{m_\text{KCl}+m_{\text{H}_2\text{O}}}100\%=1.6\%."

The percent by volume of 20 mL of alcohol diluted in water to produce 250mL solution is


"\phi=\frac{V_{al}}{V_{sol}}100\%=\frac{20}{250}100\%=8\%."